OpenStudy (anonymous):

The area of a rectangular field is 875 m2. Two adjacent sides of the field are fenced with wood costing $5/m. The remaining two sides are fenced with steel costing $10/m. The total cost of the fencing is $900. What are the dimensions of the field?

OpenStudy (anonymous):

Let the dimensions be A x B (or a x b). ab = 875 15a + 15b = 900 [ from 10a + 5a + 10b + 5b = 900 ] a + b = 60 a (60 - a) = 875 60a - a^2 = 875 a^2 - 60a + 875 = 0 Dimensions are 35 x 25 = 875 15 * (35 + 25) = 15 * 60 = 900

OpenStudy (anonymous):

Thank you so much!! :)

OpenStudy (anonymous):

you welcome (: